OpenStudy (anonymous):

Please explain how to do this ! Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3. y = -3x - 13 y = -3x + 11 y = -3x + 13 y = -3x + 1

OpenStudy (ghazi):

first assume a line y=mx +c where m is slope and c is the constant now y= -3x+c is the line pass it through the given points to find c that is (2,-5) -5=-3*2+c -5=-6+c c=1 therefore your equation is y=-3x+1

OpenStudy (anonymous):

thanks but naw can you type that in english? o.0

OpenStudy (ghazi):

thats already in english lol

OpenStudy (anonymous):

lol ok so then can u dumb it down like ALOT? =D

OpenStudy (ghazi):

well i wrote the whole algorithm of your question, that is how to find the equation of line and what do you mean by ALOT lol i dont know

OpenStudy (anonymous):

im not smart at math like at all so i need simple terms

OpenStudy (ghazi):

cool, i will go step by step, you could have written this before lol

OpenStudy (anonymous):

kk n like i said i not smart xD

OpenStudy (ghazi):

first of all consider an equation of line that is y=m*x+c where m is slpe of a line c is the constant

OpenStudy (anonymous):

k

OpenStudy (ghazi):

then put the value of slope in the assumed equation your line would be y=-3x+c

OpenStudy (anonymous):

k THANK U HERO xD lol easy terms xD

OpenStudy (ghazi):

now to find c think of the point from where your line passes through that is x=2 and y=-5 put these value in y=-3x+c -5=-3*2+c now find c and here is your answer

OpenStudy (anonymous):

kk thanks

OpenStudy (ghazi):

:D